# Large Language Model-driven Meta-structure Discovery in Heterogeneous Information Network

### v1

@@ Abstract.
Heterogeneous information networks (HIN) have gained increasing popularity in the recent decade for their ability to capture the complex relations between nodes of diverse types. Meta-structure was proposed to identify the important patterns of relations on HIN, which has been proven effective for extracting rich semantic information from HIN and facilitating graph neural networks to learn expressive representations. However, hand-crafted meta-structures pose challenges for scaling up to large and complex HINs, which draw wide research attention for developing automatic meta-structure search algorithms. Previous efforts predominantly concentrate on searching for meta-structures that have good empirical prediction performance, overlooking the explainability of discovered meta-structures. As a result, they often produce complex meta-structures that are prone to overfitting and incomprehensible to human researchers. To address this challenge, we draw inspiration from the emergent reasoning abilities of large language models (LLMs). In this paper, we propose a novel reasoning meta-structure search (ReStruct) framework that integrates LLM reasoning into the evolutionary procedure for meta-structure search. ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate the semantically feasible meta-structures. Besides, ReStruct also employs performance-oriented evolutionary operations. These two competing forces allow ReStruct to jointly optimize for semantic explainability and empirical performance of meta-structures. Finally, we also design a differential LLM explainer that can produce natural language explanations for the discovered meta-structures, and refine the explanation by reasoning through the search history. Experiments on five representative HIN datasets demonstrate that ReStruct achieve state-of-the-art performance in both node classification and link recommendation tasks. Additionally, a survey study involving $73$ graduate students shows that the meta-structures and natural language explanations generated by ReStruct are substantially more comprehensible.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Heterogeneous information networks (HINs) have been proven effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016). As a result, HINs are increasingly adopted in a wide range of applications, including social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020). To fully exploit the rich semantic information encoded in HINs, researchers proposed to use hand-crafted templates of relation sequences to model the complex proximities on HINs (Sun et al., 2011), which is known as meta-path. It was later extended to meta-structure to capture more general interaction patterns than linear paths. The recent advancement of heterogeneous GNNs showed that meta-structures can be leveraged to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019), which are crucial for various downstream applications. However, the previous works on HINs modeling predominated leverage hand-crafted meta-structures, which heavily rely on the knowledge of domain experts. Therefore, they are difficult to scale up to larger and more complex HINs that are vital for real-world applications.
@@
Motivated by the importance of domain adaptive meta-structure design, numerous recent research efforts are dedicated to designing algorithms for automatic meta-structure design. Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search model (Ding et al., 2021) to automatically identify meta-structures that could lead to superior performance in heterogeneous GNNs. However, these previous attempts solely focus on the prediction performance of meta-structures, which often lead to highly complex structures that are difficult to explain and susceptible to overfitting. Such “meta-structures” deviate from the original inspiration of meta-structures research that aims to extract semantically clear features from HINs (Huang et al., 2016).
@@
The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to address the challenges of meta-structure discovery. The scaled-up versions of LLMs show emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a). For example, researchers find chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b). Such general-purpose reasoning capability holds huge potential to make sense of the rich semantic information in HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms. That is, instead of searching for best-performing meta-structures, the reasoning power of LLM could facilitate the joint optimization for semantic explainability by emulating human reasoners. Such reasoning process could effectively prevent the overfitting of identified meta-structures, and produce human understandable knowledge from given HINs.
@@
In this paper, we propose a novel ReStruct (short for Reasoning meta-Structure search) framework that integrates LLM reasoning into an evolutionary procedure for meta-structure search. Specifically, ReStruct employs a grammar translator to encode meta-structures with nested clauses in natural language (see Figure 2). It ensures the semantic meaning of meta-structures can be readily comprehended by LLMs as natural language sentences. Besides, we define a set of basic operations that can explore the feasible adjacent meta-structures of a given meta-structure, allowing ReStruct to search the adjacent possibilities of promising candidates. Different from pure performance-oriented search, ReStruct jointly evaluates the semantic feasibility and empirical performance to determine promising candidates. On one hand, ReStruct employs a few-shot LLM predictor to evaluate the semantic similarities between given meta-structure candidates and the searched meta-structures in past experience, and use a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities. On the other hand, after the semantically feasible candidates are empirically evaluated with heterogeneous GNNs, ReStruct uses a evolutionary updater adopting the classic elimination-reproduction procedure to mimic natural selection processes that refine meta-structure candidates based on their performances. Finally, we design a differential LLM explainer that generates natural language explanations for the discovered meta-structure. It employs a chain-of-thought prompting technique to perform structural comprehension and performance attribution step-by-step. Such a reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent, not chosen ones.
@@
We evaluate ReStruct on five representative HIN datasets. Experiments show that ReStruct achieve state-of-the-art performance on both node classification and link recommendation tasks. Besides, ReStruct can generate natural language explanations as it searches through the solution space. To effectively evaluate the explainability of discovered meta-structures, we perform a user survey on $73$ graduate students with domain knowledge of HIN research. Survey results show that the meta-structure discovered by our ReStruct method is chosen as the most comprehensible one by 46.6% participants against 3 strong baselines, outperforming the second best baseline by 61.8%. More importantly, the natural language explanations generated by our method are preferred by most human participants (77.6% on average) in a head-to-head comparison with baseline methods.
@@
  • We propose a novel ReStruct framework that integrates LLM reasoning into an evolutionary meta-structure search procedure. ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, which coordinates the competing forces of empirical performance-oriented evolutionary updater and semantic similarity-oriented LLM selector. ReStruct represents a path-breaking development in meta-structure search algorithms, which could generate meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
  • We design a grammar translator to encode meta-structures as natural language sentences, which unleashes the reasoning power of LLMs to make sense of the rich semantic information on HINs. On top of this, we design a differential LLM explainer that can generate human comprehensible natural language explanations for discovered meta-structures.
  • Extensive experiments show that ReStruct achieves state-of-the-art performance on five representative datasets. More importantly, a user survey validates that ReStruct substantially outperforms baseline methods in terms of the explainability of discovered meta-structures.
@@ 2. Preliminaries
Definition 3.3. Meta-structure (Huang et al., 2016). Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures between the source node type and the target node type.
@@ 3. Methods
In this section, we introduce our proposed methods in detail. In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLM’s understanding of its semantic meanings. In Section 3.2, we design three basic operations for generating candidate meta-structures. In Section 3.3, we design two LLM modules to evaluate and select candidate meta-structures with semantic similarity orientation. In Section 3.4, we combine LLM-guided optimization with evolutionary processes to form an effective derivative-free optimization framework. Figure 1 shows the overview of our framework.
@@ 3.1. Natural Language Encoding of Meta-Structures
To facilitate the LLM’s comprehension, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2. For any meta-structure, we first traverse the structure to find all possible simple paths from the source node to the target node. Each resulting path is equivalent to a meta-path connecting the source and the target, decomposed from the original meta-structure. For each path, we encode it into a natural language sentence using nested clauses signified by THAT, which is a commonly-used grammar in English and thus hypothesized to be well-comprehended by the LLM. After obtaining the natural language encodings of the decomposed meta-path, i.e., sub-logics, we further combine them using a conjunction word AND to indicate the logical summation effect.
@@ 3.2. Basic Operations for Candidate Meta-Structure Generation
To generate comprehensive candidates for LLM selection and meanwhile ensure validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing Lego. Examples are shown in Figure 3.
  • INSERTION. This operation replaces one edge of the original meta-structure with a component.
  • GRAFTING. This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively.
  • DELETION. This operation deletes a certain amount of nodes from the original meta-structure, and tries connecting all neighbors back.
In our experiments, we take all meta-paths with no more than $2$ nodes as components for INSERTION, and all meta-paths with no more than $3$ nodes as components for GRAFTING. DELETION does not require components as input, as it regards all existing nodes on the original meta-structure as operation candidates. Note that using components with more nodes can bring larger exploration space at each step, but may also result in chaos and confusion for the LLM, and thus we leave it as future work to investigate the optimal component settings.
@@
3.3. Semantic Similarity-oriented LLM modules for Candidate Selection
To regulate the explainability of discovered meta-structures, we design two LLM modules that evaluate and select candidate meta-structures in a semantic similarity-oriented manner. We illustrate the interaction processes between the main program and the LLM agents in Figure 4.
@@ 3.3.1. Few-shot LLM Predictor
Following the execution of the aforementioned basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates. This collection comprises meta-structures that have been previously encountered and evaluated in earlier generations, as well as entirely novel ones. To inform the decision-making process with insights gleaned from previous evaluations, we design a few-shot LLM predictor to predict the performance $\hat{p}$ of each candidate based on a performance pool that stores the structures and corresponding performances of seen meta-structures. Additionally, the few-shot LLM predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate. Intuitively, if the few-shot LLM predictor deems a candidate highly similar to a counterpart in the performance pool, it is inclined to assign higher confidence to this prediction. This is grounded in the notion that structural similarity is likely indicative of functional similarity. An illustrative prompt-response round is exemplified in Step 1 of Figure 4.
@@ 3.3.2. Similarity-oriented LLM Selector
Receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design an Similarity-oriented LLM selector to make the final decision of picking one single candidate to enter the next generation. During this process, the Similarity-oriented LLM selector is expected to jointly consider, and perhaps trade-off between, a meta-structure’s semantic meanings, the expected outcomes of choosing it, as well as the credibility of this outcome expectation. An illustrative prompt-response round is exemplified in Step 2 of Figure 4.
@@ 3.4. Performance-oriented Evolutionary Updater
With closed-source LLM modules in the loop, it is not feasible for us to obtain the gradient for optimizing meta-structure search. Therefore, we operationalize a derivative-free optimization framework with inspirations from the genetic algorithm. Specifically, we maintain a population of $N$ individuals, each representing a distinct meta-structure. In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task such as recommendation and node classification. After evaluation, the underperforming meta-structures are eliminated from the population. The surviving meta-structures undergo a reproduction phase, where duplication occurs with probabilities proportional to their performances. In essence, this phase uses promising meta-structures to replenish the population to its original size. Both elimination and reproduction processes mirror natural selection mechanisms that enable species evolution in the wild. After getting the modified population, we feed it into the LLM to improve each individual meta-structure. Under the evolutionary framework, this can be regarded as a way of targeted mutation, as in this process new nodes and/or edges can be generated and some of the existing nodes and/or edges will be removed. The modified population will be further evaluated at the onset of the next generation, forming a loop of derivative-free optimization.
3.5. Differential LLM Explainer

### v2

@@ Abstract.
Heterogeneous information networks (HIN) have gained increasing popularity in recent years for capturing complex relations between diverse types of nodes. Meta-structures are proposed as a useful tool to identify the important patterns in HINs, but hand-crafted meta-structures pose significant challenges for scaling up, drawing wide research attention towards developing automatic search algorithms. Previous efforts primarily focused on searching for meta-structures with good empirical performance, overlooking the importance of human comprehensibility and generalizability. To address this challenge, we draw inspiration from the emergent reasoning abilities of large language models (LLMs). We propose ReStruct, a meta-structure search framework that integrates LLM reasoning into the evolutionary procedure. ReStruct uses a grammar translator to encode the meta-structures into natural language sentences, and leverages the reasoning power of LLMs to evaluate their semantic feasibility. Besides, ReStruct also employs performance-oriented evolutionary operations. These two competing forces allow ReStruct to jointly optimize the semantic explainability and empirical performance of meta-structures. Furthermore, ReStruct contains a differential LLM explainer to generate and refine natural language explanations for the discovered meta-structures by reasoning through the search history. Experiments on eight representative HIN datasets demonstrate that ReStruct achieves state-of-the-art performance in both recommendation and node classification tasks. Moreover, a survey study involving $73$ graduate students shows that the discovered meta-structures and generated explanations by ReStruct are substantially more comprehensible. Our code and questionnaire are available at https://github.com/LinChen-65/ReStruct.
@@
copyright: acmcopyrightjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671965isbn: 979-8-4007-0490-1/24/08conference: the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–28, 2024; Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Social networksccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Heterogeneous information networks (HINs) are effective in jointly modeling network topology and multi-typed relations (Shi et al., 2016), leading to their widespread adoption across various applications, such as social media (Xu et al., 2019), information retrieval (Wang et al., 2019), and recommender systems (Han et al., 2020; Ding et al., 2024). To fully exploit the rich semantic information encoded in HINs, researchers have proposed to use meta-paths, which are templates of relation sequences to model the complex proximity on HINs (Sun et al., 2011). They were later extended to meta-structures to capture more general interaction patterns beyond linear paths (Huang et al., 2016). These meta-structures have been successfully utilized in heterogeneous graph neural networks (GNNs) to learn expressive representations for HINs (Xu et al., 2019; Wang et al., 2019) for completing downstream tasks. However, the reliance on hand-crafted meta-structures, which depend on domain experts’ knowledge, makes it challenging to scale up to larger and more complex HINs that are commonly encountered for real-world applications.
@@
Driven by the importance of domain adaptation, recent research efforts have been dedicated to developing algorithms for automatic meta-structure search. Researchers propose to use genetic algorithm (Han et al., 2020), deep reinforcement learning (Peng et al., 2021) and differentiable neural architectural search models (Ding et al., 2021) to automatically identify meta-structures that can enhance the performance of heterogeneous GNNs. However, these previous attempts primarily focus on the prediction performance of meta-structures, often resulting in highly complex structures that are challenging to interpret and prone to overfitting. Such “meta-structures” deviate from the original inspiration of meta-structure research that aims to extract semantically clear features from HINs (Huang et al., 2016).
@@
The recent breakthrough in large language models (LLMs) (Floridi and Chiriatti, 2020) offers a unique opportunity to tackle the challenges of meta-structure discovery. The scaled-up versions of LLMs have exhibited emergent abilities for a wide range of complex tasks that go beyond auto-regression token generation (Wei et al., 2022a). For example, researchers have found that chain-of-thoughts prompting can effectively unlock LLMs’ reasoning capability for commonsense, mathematical, and logical problems (Wei et al., 2022b). Such a general-purpose reasoning capability holds huge potential for comprehending the rich semantic information and produce human understandable knowledge from given HINs, which could be path-breaking to current performance-oriented meta-structure search algorithms.
@@
In this paper, we propose a novel framework named ReStruct (short for Reasoning meta-Structure search) that integrates LLM reasoning into an evolutionary procedure for meta-structure search. In this framework, we design a grammar translator to encode meta-structures into natural language sentences with nested clauses (see Figure 2), ensuring that their semantic meanings can be readily comprehended by LLMs. Besides, we define a set of basic operations to modify a given meta-structure, allowing ReStruct to explore its adjacent possibilities in a valid space. Unlike pure performance-oriented search, we anticipate ReStruct to evaluate both semantic feasibility and empirical performance to identify promising candidates. To this end, we first design a few-shot LLM predictor to estimate the performance of meta-structure candidates with access to previously evaluated meta-structures from a history pool, followed by a similarity-oriented LLM selector to identify the most promising candidates based on the semantic similarities. After empirically evaluating the chosen candidates with heterogeneous GNNs, we design an evolutionary updater adopting the classic elimination-reproduction procedure to refine meta-structure candidates based on their performances. Finally, we design a differential LLM explainer that generates natural language explanations for the discovered meta-structure. It employs a chain-of-thought prompting technique to perform step-by-step structural comprehension and performance attribution. This reasoning process generates high-quality explanations by explicitly comparing the chosen meta-structures and the adjacent yet unchosen ones.
@@
We evaluate ReStruct on eight representative HIN datasets. Experiments show that ReStruct achieves state-of-the-art performance on both recommendation and node classification tasks, and generates meaningful explanations as it searches through the solution space. To effectively assess the explainability of the discovered meta-structures, we conduct a user survey on $73$ graduate students with domain knowledge in HIN research. According to the survey results, 46.6% of the participants consider the meta-structure discovered by ReStruct as the most comprehensible compared with three strong baselines, outperforming the second best baseline by 61.8%. Moreover, the natural language explanations generated by ReStruct are significantly preferred by the majority (77.6% on average) in a head-to-head comparison with baseline methods.
@@
  • We propose a novel ReStruct framework that integrates LLM reasoning into an evolutionary meta-structure search procedure. ReStruct jointly optimizes the empirical prediction performance and semantic explainability of meta-structures, by coordinating the competing forces of an evolutionary updater and a semantic similarity-oriented LLM selector. This represents a significant advancement in meta-structure search algorithms, enabling the generation of meta-structures that represent human digestible knowledge on HINs and are less prone to overfitting.
  • We design a grammar translator to encode meta-structures as natural language sentences, which unleashes the reasoning power of LLMs to make sense of the rich semantic information on HINs. On top of this, we design a differential LLM explainer that can generate human-comprehensible natural language explanations for discovered meta-structures.
  • We conduct extensive experiments to reveal ReStruct’s state-of-the-art performance on eight representative datasets. Furthermore, we carry out a user survey to validate that ReStruct substantially outperforms baseline methods in terms of the comprehensibility of discovered meta-structures and usefulness of generated explanations.
@@ 2. Preliminaries
Definition 3.3. Meta-structure (Huang et al., 2016). Given an HIN $G$, a meta-structure $T$ is a generalization of the meta-path to allow for the existence of graph structures beyond linear connections between the source node type and the target node type.
@@ 3. Methods
In this section, we provide a detailed introduction to our proposed methods. In Section 3.1, we elaborate our novel design of natural language encoding of meta-structures to facilitate LLMs’ understanding of its semantic meanings. In Section 3.2, we introduce our design of three basic operations for generating candidate meta-structures. In Section 3.3, we design two LLM agents to evaluate and select candidate meta-structures with semantic similarity orientation. In Section 3.4, we combine LLM-guided optimization with evolutionary processes to form an effective derivative-free optimization framework. The overview of our framework is shown in Figure 1.
@@ 3.1. Natural Language Encoding of Meta-Structures
Previous works represent meta-structures either as matrices or sets of numbers (Han et al., 2020; Ding et al., 2021), which can be challenging to interpret in terms of their semantic meanings. As a result, this poses obstacles for LLMs to effectively comprehend such representations. To address this limitation and enhance LLMs’ comprehension of meta-structures, we design a grammar translator module to encode each meta-structure into a natural language sentence, as shown in Figure 2. For a given meta-structure, we begin by traversing its structure to find all possible simple paths connecting the source node to the target node. Each resulting path is equivalent to a meta-path decomposed from the original meta-structure. Next, we encode each path into a natural language sentence using nested clauses signified by a conjunction word THAT, which is a commonly-used grammar in English and thus expected to be well-comprehended by the LLM. In each clause, the central verb connecting two entities is the semantic meaning corresponding to the edge connecting two nodes. After obtaining the natural language encodings of the decomposed meta-paths, i.e., sub-logics, we further combine them using another conjunction word AND to convey the logical summation effect.
@@ 3.2. Basic Operations for Candidate Meta-Structure Generation
To generate comprehensive candidates for LLM selection while ensuring validity, we define three basic operations for modifying any meta-structure, and design a set of components for these operations, analogous to playing with Lego blocks. Examples are shown in Figure 3.
  • INSERTION. This operation replaces one edge of the original meta-structure with a component. It introduces new connections and expands the structure.
  • GRAFTING. This operation takes a component, finds two nodes in the original meta-structure with the same type as the chosen component’s first and last node, and merges them respectively. It creates branching structures to enhance expressiveness.
  • DELETION. This operation removes a certain amount of nodes from the original meta-structure, and reconnects the remaining nodes to ensure the structure remains valid.
In our experiments, we take all meta-paths with no more than $2$ nodes as components for INSERTION, and all meta-paths with no more than $3$ nodes as components for GRAFTING. DELETION does not require components as input, as it regards all existing nodes on the original meta-structure as operation candidates. While using components with more nodes expands the exploration space, it may also introduce complexity and confusion for the LLM. We leave it as future work to investigate the optimal component settings.
@@
[…]
Basic operations of exploring adjacent meta-structures.
3.3. Semantic Similarity-Oriented LLM Agents for Candidate Selection
To regulate the explainability of discovered meta-structures, we design two LLM agents that evaluate and select candidate meta-structures in a semantic similarity-oriented manner. We illustrate the interaction processes between the main program and the LLM agents in Figure 4.
@@ 3.3.1. Few-Shot LLM Predictor
After applying the basic operations on each meta-structure, we derive a set of one-step neighbors as potential candidates. These neighbors can include meta-structures that have been encountered and evaluated in earlier generations, as well as entirely new ones. To leverage insights from previous evaluations and guide the decision-making process, we design an LLM agent as a few-shot LLM predictor (abbreviated as “predictor” below). This predictor estimates the performance $\hat{p}$ of each candidate through instruction tuning on a small set of structure-performance pairs sampled from a performance pool that records meta-structures in all previous rounds. Additionally, the predictor is asked to provide a self-estimated confidence value $\hat{c}$ for each prediction, resulting in a $(\hat{p},\hat{c})$ pair associated with each candidate. Intuitively, if the predictor considers a candidate to be highly similar to a counterpart in the performance pool, it is likely to predict a similar performance and assign higher confidence to this prediction. This is grounded in the understanding that structural similarity often implies functional similarity. An illustrative prompt-response round is exemplified in Step 1 of Figure 4.
@@ 3.3.2. Similarity-Oriented LLM Selector
Upon receiving a set of candidates and their corresponding predicted performances from the few-shot LLM predictor, we design another LLM agent, i.e., a similarity-oriented LLM selector (abbreviated as “selector” below), to make the final decision of selecting one single candidate to proceed to the next generation. During this process, the selector is expected to consider multiple factors simultaneously, and potentially trade-off between them in order to make the optimal decision (see Appendix B,C). These factors include: (1) Semantic meanings, which reflect the relevance and alignment of the meta-structure with the desired objectives and requirements. (2) Structural complexities, which indicates the potential risks of overfitting. (3) Expected outcomes provided by the few-shot LLM predictor, which indicates the potential benefits from selecting a particular meta-structure in terms of performance improvement. (4) Credibility of outcome expectation also provided by the few-shot LLM predictor, which reflects the reliability and trustworthiness of the predictions. An illustrative prompt-response round is exemplified in Step 2 of Figure 4.
@@ 3.4. Performance-Oriented Evolutionary Updater
With closed-source LLM modules in the loop, it is not feasible for us to obtain the gradient for optimizing meta-structure search. Therefore, we operationalize a derivative-free optimization framework with inspirations from the genetic algorithm. Specifically, we maintain a population of $N$ individuals, each representing a distinct meta-structure. In every generation, we first evaluate the performance of each meta-structure by using it to train a GNN for the given downstream task. After evaluation, the underperforming meta-structures are eliminated from the population. The surviving meta-structures undergo a reproduction phase, where duplication occurs with probabilities proportional to their performances. In essence, this phase uses promising meta-structures to replenish the population to its original size. Both elimination and reproduction processes mirror natural selection mechanisms that enable species evolution in the wild. After getting the modified population, we feed it into the aforementioned LLM agents to for a new round of individual meta-structure improvement. This step can be seen as a way of targeted mutation within the evolutionary framework, as new nodes and/or edges can be generated and some of the existing nodes and/or edges may be removed. The modified population will be re-evaluated at the onset of the next generation, forming a loop of derivative-free optimization. In summary, by utilizing this evolutionary optimization framework, we can iteratively search for and improve meta-structures without relying on gradient-based optimization methods.
3.5. Differential LLM Explainer Agent